# Advances in Embodied Navigation Using Large Language Models: A Survey

### v4

@@ 3.3.3 Comparison
3.4 Other Applications of LLMs in Embodied Intelligence
1)VoxPoser [3] focuses not on embodied navigation works but on embodied AI. VoxPoser is a novel approach that utilizes LLMs to guide robots in performing various tasks without the need for predefined motion primitives. Through LLMs, it can extract crucial cues and constraints from natural language instructions to generate robot motion trajectories. For example, when given the command ”Open the topmost drawer, but be careful of the vase,” LLMs can understand the need to grasp the drawer handle and pull it outward while avoiding the vase. Subsequently, by assembling 3D value maps, robots can execute these actions in the real environment without the need for additional training.
Unlike traditional methods, VoxPoser not only addresses the issue of instruction execution but also generates robust robot trajectories in dynamic environments. It leverages the code-generation capabilities of LLMs to translate language instructions into Python code, which invokes perception APIs to obtain spatial information about objects in the environment, thus creating 3D voxel maps. These maps serve as the objective function for motion planners, guiding robots in action execution. The uniqueness of this approach lies in its ability to allow robots to perform tasks based on natural language instructions in unknown scenarios, thereby achieving zero-shot generalization to open instructions.
Overall, VoxPoser provides a new paradigm for robots to perform complex tasks, making full use of the capabilities of language models to interpret and execute natural language instructions. Its application potential extends beyond robot operation to other domains such as autonomous driving vehicles or smart home systems.
2)ALFRED [82] (Action Learning From Realistic Environments and Directives) serves as both a benchmark and dataset for the training and evaluation of models across a spectrum of embodied intelligence tasks. The central focus of this benchmark is to examine the efficacy of AI agents in carrying out domestic tasks, based on detailed natural language directives. The dataset comprises 25,743 English instructions that correspond to 8,055 expert demonstrations. Each demonstration encompasses approximately 50 steps, culminating in a total of 428,322 paired image-action sequences.
ALFRED tasks are highly complex, often necessitating that agents execute a chain of sub-tasks involving irreversible state alterations, such as breaking an object or transferring food between containers. Included in the dataset are high-level objectives like ”rinse a mug and place it in the coffee maker”, as well as more granular instructions such as ”walk to the coffee maker on the right”. Robots interacting with human habitats must learn to associate natural language with their surroundings. ALFRED outperforms existing vision and language task datasets by offering a wealth of expert demonstrations in simulation environments that are more intricate in terms of sequence length, action space, and language complexity.
Contrary to the advancements made possible by datasets like MP3D, these resources often fail to address the complexities arising from task-oriented object interactions. ALFRED mandates that agents extract essential information from natural language directives, encompassing both the goals and potential execution steps for tasks. Agents are also required to deconstruct high-level tasks into executable actions or sub-tasks and plan an optimal path for task completion. Additionally, they must identify various objects and environmental features from a first-person perspective and engage in complex interactions with environmental objects, such as picking up items, manipulating switches, or navigating to specific locations.
By offering a visually and physically realistic simulated environment that enables translation from language to action sequences and interactions, ALFRED tackles numerous challenges evident in real-world scenarios. This includes translating human language into robotic actions for household tasks. Models capable of overcoming these challenges will begin to narrow the divide between real-world, language-driven robots and their simulated analogs.
3)PaLM-E [83] (Pre-training across Language and Modality Embodied) serves as a multimodal language model proficient in integrating visual, linguistic, and sensor modalities to facilitate more grounded reasoning for complex real-world challenges. Specializing in robotic planning, visual question-answering, image description, and linguistic tasks, PaLM-E exhibits robust transfer learning capabilities. It trains as a generalized, efficient embodied reasoner by merging embeddings of real-world entities in its dataset, making it skillful in multi-agent decision-making and transfer learning. Experiments indicate that concurrent training on diverse robotic tasks and datasets bolsters its performance and generalization capacity. Moreover, PaLM-E excels in language-conditioned policies and self-supervised learning, enabling both zero-shot and one-shot generalizations.
While LLMs display substantial reasoning capabilities across an array of domains, they are constrained by a grounding issue. Previous research, such as work by Ahn et al. [84], combined the outputs of LLMs with learned robotic policies and affordance functions. However, these efforts were limited as they supplied only textual inputs to the LLM, insufficient for tasks demanding geometric scene understanding. Training LLMs on expansive textual data may yield representations pertaining to the physical world, yet integrating these with real-world visual and sensor modalities is essential for addressing broader, grounded challenges in computer vision and robotics.
PaLM-E augments this by directly assimilating continuous sensor modalities within the agent, thereby enriching the language model’s capacity for informed, sequential decision-making in real-world contexts. Inputs like images and state estimates are encoded in the same latent space as linguistic tokens, and are processed through the self-attention layers of a Transformer-based LLM analogously to text. Driess et al. initiated this process with a pre-trained LLM, incorporating continuous inputs via encoders. These encoders undergo end-to-end training to produce sequential decisions in natural language, interpretable by agents through fine-tuned, low-level policies or by providing specific query responses.
Utilizing the Transformer architecture and drawing from a diverse set of training data that encompasses text, images, and audio, PaLM-E can comprehend and manage multiple information types. It boasts zero-shot learning capabilities, enabling it to tackle previously unencountered tasks via inference and generation. The model accommodates a broad array of inputs like text, images, and audio, rendering it versatile for complex tasks. Its generative prowess extends to the production of natural language text, image descriptions, and beyond, thereby making it applicable across various natural language generation domains.
4)RT-2 [85] (Robotics Transformer 2) serves as an avant-garde model crafted to adapt web-sourced visual and linguistic insights into robotic control functions. Composed of dual core elements—a pre-trained vision-language model and a reinforcement learning-based robotic control policy—RT-2 leverages expansive web data to cultivate comprehensive visual and linguistic awareness. Exhibiting remarkable proficiency in both simulated and authentic environments, the model excels in executing intricate operations such as object manipulation and door-opening tasks.
Contemporary researchers like Brohan et al. contend that LLMs function primarily at a macroscopic level, essentially acting as state machines that dictate orders to independent, microscopic controllers. These low-level controllers often lack access to the rich semantic knowledge harbored within web-scale models during their developmental stages. They probe the feasibility of incorporating pre-trained, large-scale Visual Language Models directly into these microscopic controllers to amplify their generalization and emergent semantic reasoning faculties. In response, they develop visual-language models tailored for open-vocabulary visual question-answering and dialogue, programmed to generate low-level robotic actions, while concurrently addressing broader web-scale visual-language challenges.
Through the refinement of existing models such as PaLM-E [83] and PaLI-X [86] into RT-2-PaLM-E and RT-2-PaLI-X, Brohan et al. engineer high-efficacy robotic schemes that inherit the advanced generalization and emergency response features facilitated by web-scale visual-language pre-training. This unambiguous yet multifaceted strategy elucidates the immense potential for robotic technologies to gain directly from the strides made in visual-language models, thereby aligning the progression of robotics with concurrent advancements in complementary fields.
Nonetheless, deploying large-scale Visual-Language-Action (VLA) models for real-time operations necessitates considerable computational resources, thereby raising issues for scenarios that demand high-frequency control. A promising avenue for ensuing research encompasses the exploration of quantization and distillation techniques, which could facilitate the functioning of these models at elevated speeds or on more economically viable hardware platforms.
3.5 State-of-the-Art Datasets
@@
1)MatterPort3D [87] (MP3D) is a comprehensive dataset, predominantly comprising RGB-D scenes. It features 10,800 panoramic views generated from 194,400 RGB-D images and spans 90 expansive architectural settings, primarily focusing on 3D indoor environments. The dataset encompasses various data modalities such as RGB images, depth images, and semantic annotations. Annotations are supplied for facets like scene surfaces, camera poses, and semantic segmentation. Given its voluminous collection of scenes and precise global perspectives, MP3D serves as a pivotal resource for tasks reliant on self-supervision or alternative methods.
As a cornerstone in the field, MP3D has influenced the development of numerous subsequent datasets, which have either adapted or expanded upon its structure. This renders it an invaluable asset for simulating Embodied AI in computational settings. Nevertheless, the dataset’s primary focus on indoor residential environments circumscribes its applicability. Unique to MP3D, each viewpoint incorporates a 360-degree panoramic image, encompassing both depth and color information. The dataset meticulously samples viewpoints at human height throughout these spaces, ensuring camera poses are aligned. Data are amassed across a range of applications, from semantic segmentation to object and regional categorization.
For each object surface, MP3D furnishes multiple views captured from diverse angles and distances, thereby enabling the learning and prediction of surface properties contingent on these perspectives. The dataset places a significant emphasis on visual perception and scene understanding, making it highly versatile for a plethora of computer vision tasks including object detection, scene segmentation, and 3D reconstruction. For Embodied AI research, MP3D stands as an exemplary dataset, offering a data-rich environment that enhances agent-environm-ent interactions.
2)TOUCHDOWN [88] is a specialized dataset designed for navigation and spatial reasoning, utilizing Google Street View to construct an expansive outdoor environment. The dataset encompasses 39,641 panoramic images and 61,319 edges, all originating from New York City. Each viewpoint in the dataset features a 360-degree RGB panorama. To compile this dataset, TOUCHDOWN employs a ”treasure hunt” methodology, involving two roles: the writer and the finder. The writer crafts instructions to enable the finder to navigate toward and locate a specific target, without prior knowledge of the correct route or destination. Upon task completion, the finder documents the followed path and elucidates their navigation choices.
TOUCHDOWN consists of two primary tasks: Navigation and Spatial Description Resolution (SDR). The Navigation task emphasizes spatial reasoning, with instructions focusing on the spatial relationships between the agent and its environment. The SDR task, on the other hand, demands linguistic expertise to decipher relationships among objects and their spatial configurations. Each of these tasks presents its own set of unique challenges.
Contrary to MP3D [87], which is confined to indoor scenarios, TOUCHDOWN predominantly features outdoor settings. This makes it an ideal dataset for training and evaluating models in tasks that require outdoor navigation and spatial understanding.
3)Room-to-Room [89](R2R) is a specialized dataset designed for VLN, extending the data from MP3D to include a substantial corpus of natural language data pertinent to navigation. R2R characterizes a series of navigational paths within these 3D indoor settings, each consisting of a sequence of specific locations and orientations, termed ’viewpoints.’ Accompanying each path are one or more natural language instructions that guide navigation from the origin to the destination.
Unlike MP3D, which is more general-purpose, R2R employs these indoor settings for highly specialized navigation tasks, necessitating a nuanced understanding of natural language and complex decision-making. Navigating intricate indoor terrains often demands capabilities in long-term memory and reasoning. A model utilizing R2R not only needs to comprehend natural language directives but must also correlate them with the visual landscape. Upon this synthesis, the model is required to strategize an efficient route from the starting location to the specified endpoint. Through these multiple dimensions, R2R offers an intricate and demanding research platform for those specializing in vision-and-language navigation.
4)CVDN [90] is an extensive dataset comprising over 2,000 tailored ”human-to-human” dialogues, set in a simulated, photorealistic domestic environment. Developed as an extension of the R2R framework and formulated in English, CVDN aims to cultivate navigation agents capable of functioning in residential and commercial spaces via linguistic directives. Notably, these agents have the ability to pose focused questions during instances of uncertainty regarding subsequent navigational steps.
Furthermore, CVDN can be leveraged to develop agents endowed with specialized environmental knowledge, thereby facilitating automated linguistic guidance for individuals in unfamiliar contexts—such as direction-seeking within an office complex. The high fidelity of the simulated environment ensures that agents trained on CVDN can generalize skills to real-world settings.
Diverging from R2R, dialogues within CVDN typically contain nearly triple the word count and describe paths averaging over thrice the length. Thus, CVDN provides a significant avenue for the exploration of in-situ navigational interactions and holds potential for the training of agents proficient in both navigation and dynamic inquiry within human-centric environments.
5)REVERIE [91] is a comprehensive dataset featuring 10,567 panoramic images spread across 90 buildings and encompassing 4,140 target objects. Accompanying these are 21,702 crowd-sourced instructions, averaging 18 words each. While it shares visual data attributes like RGB images with earlier datasets, REVERIE also incorporates enhanced semantic particulars related to the Goal Object, such as color, along with natural language directives.
Distinct from conventional VLN datasets, REVERIE includes a myriad of complex and diverse indoor environments. Moreover, its instruction set diverges from the step-by-step commands typical of R2R datasets to offer semantic-level directives more congruent with real-world tasks and human-agent interactions. For example, instead of sequential directives like ’Start at the entrance and proceed to the red sofa,’ REVERIE provides semantically rich tasks such as ’Fold the towel in the bathroom with the fishing theme.’
This semantic depth, however, introduces a layer of complexity due to potential ambiguities (e.g., ’Fold the towel in the bathroom with the fishing theme’ could signify either ’Fold the fishing-themed towel’ or ’in the fishing-themed bathroom’). Consequently, agents trained on the REVERIE dataset require enhanced long-term memory, semantic comprehension, and environmental reasoning abilities. Such attributes make REVERIE particularly suited for remote entity referencing and multi-task learning endeavors. In summary, REVERIE offers a multifaceted, challenging platform conducive to advancements in VLN and multimodal learning.
6)RxR [92](Room-Across-Room) is an expansive dataset tailored for Multilingual VLN tasks. As an extension of the Room-to-Room (R2R) dataset, it inherits substantial visual and navigational data, thereby enriching the research landscape with multilingual capabilities. The dataset features natural language navigation instructions in multiple languages, including but not limited to, English, German, and Spanish. This linguistic variety significantly broadens the scope for multilingual VLN endeavors.
Furthermore, RxR introduces dense spatiotemporal grounding, which entails comprehensive action execution and natural language interpretation directives, calibrated to specific temporal intervals and spatial coordinates within the navigation trajectory. Similar to REVERIE, RxR augments the dataset with additional layers of semantic data pertaining to target objects and multiple languages.
This compounded complexity poses new sets of challenges: the navigation agents must master multilingual understanding, spatiotemporal reasoning, and long-term planning. Owing to its linguistic diversity, RxR serves as a multifaceted, demanding research substrate, suitable for advancing the state-of-the-art in multilingual VLN.
7)SOON [93], also known as the FAO dataset, is a language-guided navigation dataset grounded in 3D housing environments featuring authentic image panoramas. The dataset encompasses 90 distinct residential settings, each paired with a natural language instruction that delineates a target location. Unique to SOON, the construction of these instructions is organized around five subtasks: delineation of target attributes, explication of relationships between the target and neighboring objects, exploration of the target’s regional setting, scrutiny and description of adjacent areas, and the subsequent condensation of these elements into a three-sentence narrative.
In addition to offering an extensive array of 3,848 instructions characterized by a diverse vocabulary of 1,649 words, the SOON dataset furnishes high-fidelity annotations, which include 3D environmental models, RGB-D imagery, semantic segmentation, and depth data. Remarkably, the average instruction length in SOON stands at 38.6 words, surpassing the 26.3-word and 18.3-word averages observed in REVERIE and R2R, respectively. Most instructions oscillate between 20 and 60 words, affording robust descriptive potential.
In contrast with other datasets, SOON stands out for its elevated linguistic intricacy and ecological diversity, coupled with a rich, realistic representation of environmental contexts. Consequently, the SOON dataset provides an exhaustive and reliable substrate for advancing research in the field of vision and language-guided navigation.
8)ProcTHOR [66], also known as PROCTHOR-10K, is a dataset comprising 10,000 algorithmically generated 3D residential scenes specifically tailored for the training and assessment of embodied AI agents. The dataset is a byproduct of the PROCTHOR framework, which is designed to facilitate the large-scale generation of diverse, interactive, and customizable virtual environments. Each scene within the dataset is distinct, furnished with a multiplicity of objects, textural variations, and illumination conditions, thereby constituting a rich and complex milieu for embodied agents to operate, interact, and manipulate.
Intended to support an array of embodied AI tasks—ranging from navigation and object manipulation to comprehensive scene understanding—PROCTHOR-10K has been employed for the training and evaluation of cutting-edge embodied AI models. The dataset has demonstrated robust performance across various industry benchmarks, including but not limited to Habitat 2022, AI2-THOR Rearrangement 2022, and RoboTHOR challenges.
9)R3ED [94] is a dataset encompassing over 5,800 point clouds and approximately 22,400 annotated 3D bounding boxes, sourced from seven distinct real-world indoor settings via a dense sampling technique. This dataset aspires to address the limitations inherent in previous embodied AI datasets, which predominantly rely on synthetic data to mimic indoor landscapes and robotic motions. While synthetic data offers convenience, it tends to undermine real-world performance.
Conversely, R3ED offers authentic point cloud data following each agent’s move, furnishing researchers with the opportunity to examine active vision through a more realistic and standardized lens. The point cloud data is meticulously annotated with 3D bounding boxes that categorize 12 types of objects, inclusive of prevalent furniture and appliances. Furthermore, the geometrical relationships among these point clouds are explicitly annotated using a specialized storage path format. To augment the dataset’s utility, the authors have introduced a 3D divergency policy aimed at directing robotic movements for data collection, thereby enhancing visual performance in unfamiliar environments.
@@
Table 6: Comparison of Datasets. This table is a comparison among nine datasets, including seven evaluation methods.
@@
10)X-Embodiment[95] is an embodied robot manipulation dataset, comprises over one million robot trajectories from 22 disparate robot instances, illustrating 527 skills across 160,266 tasks. The dataset amalgamates data from 60 pre-existing robot datasets, sourced from 34 international robotics research labs, and standardizes them into a unified RLDS data format[96]. This format efficiently serializes record files and accommodates a myriad of action spaces and input modalities, from varying numbers of RGB cameras to depth cameras and point clouds, thereby facilitating streamlined data loading across all primary deep learning frameworks. Transformer-based policies trained on this dataset have shown marked positive transfer capabilities across different robotic instances.
Recent developments in machine learning and artificial intelligence indicate that large-scale, diversified datasets are pivotal for the creation of competent AI systems. Such is evident in domains like natural language processing and computer vision, where large, high-capacity models trained on diverse data have effectively addressed a range of downstream applications. This trend has catalyzed the widespread adoption of pretrained models as foundational assets across applications. In the realm of robotics, models have traditionally been specialized for specific applications, robots, and environments. However, there is burgeoning interest in the development of ”generalist” robotic policies capable of efficiently adapting to novel robots, tasks, and settings. Large-scale, multipurpose models, often trained on diverse datasets, consistently outperform task-specific counterparts trained on more limited datasets.
Prompted by the efficacy of pretraining expansive vision or language models on varied data, DeepMind posits that the construction of universally applicable robotic policies necessitates X-embodiment training. This approach utilizes data aggregated from multiple robotic platforms. Although individual robotic learning datasets might be too specialized, their combination provides a more comprehensive representation of environmental and robotic variations. The development of methodologies to optimally leverage this X-embodiment data from diverse labs, robots, and settings is indispensable for the creation of universally applicable robotic policies. While current datasets might not yet match the generalization capabilities exhibited by large language models, the collaborative utilization of such resources has the prospective capacity to achieve comparable coverage. Consequently, DeepMind contends that research into X-embodiment robotic learning is of paramount importance at this stage.
In summary, this section provides an overview of ten seminal datasets, delineating their structures and merits. Table 6 offers a comparative analysis among these datasets, with more exhaustive information available in the References. These datasets provide diverse and rich resources for LLM research in the navigation domain, facilitating the optimization and development of models across various environments and tasks. MP3D offers high-quality 3D indoor environment data, aiding LLMs in understanding complex indoor scenes; TOUCHDOWN provides urban streetscape data suitable for studying natural language navigation in outdoor environments; R2R features real user-generated navigation instructions, enhancing the naturalness and diversity of the data; RXR supports multi-language navigation instructions, improving the model’s cross-language adaptability; REVERIE provides target-oriented navigation task data, promoting fine-grained navigation research; ProcTHOR offers synthetic virtual environments, supporting large-scale model training; R3ED introduces memory modules, supporting long-term navigation tasks; CVDN includes human-to-human dialogues, guiding agents in environmental navigation, suitable for studying human-machine interaction and collaborative navigation. By integrating high-quality data and multimodal inputs, these datasets advance LLMs in natural language processing, visual understanding, and multi-task learning, enhancing the intelligence and practicality of navigation systems. Broadly, both contemporary and forthcoming datasets are gravitating towards the inclusion of real-world 3D data, multi-modal inputs, higher-order instructions, and provisions for human oversight and direction.
4 Challenges and Future Directions
LLMs also face many challenges and limitations in navigation applications. In terms of datasets, each dataset has its unique strengths and weaknesses for different types of navigation and language guidance tasks. Matterport3D, R2R, RXR and REVERIE [87, 89, 92, 91] are suitable for indoor navigation tasks due to their high realism and rich annotations; TOUCHDOWN [88] is suitable for navigating urban environments; ProcTHOR [66] is applicable to the need for controlled and scalable environments; and R3ED [94] is suitable for long-term navigation tasks by introducing a memory module. Taking several datasets as examples, the MP3D dataset comes from real-world indoor environments, covering a wide range of indoor scenes such as houses and offices, with high diversity and containing detailed object labeling and semantic information, which can provide high-quality RGB-D images and 3D reconstruction data, but it also increases the difficulty of model training and inference due to the complexity of the scenes, and requires a high hardware configuration  [87]. R2R and REVERIE extend on the basis of MP3D. R2R’s scene is based on MP3D while providing real user-generated navigation instructions, but the number is limited and not diversified enough; REVERIE not only provides instructions, but also includes descriptions of the target objects, which is suitable for more fine-grained and goal-oriented navigation tasks, but with higher task complexity, and higher requirements for processing and storage. Higher. In contrast to the previous indoor datasets, TOUCHDOWN provides urban street scene data with natural language descriptions to guide navigation tasks, which is suitable for outdoor tasks, but urban scenes are more complex and the annotation quality of natural language descriptions may be inconsistent.ProcTHOR provides synthetic virtual environments, which allow the user to customize the complexity and layout of the environment, and therefore can generate a large number of scenes. CVDN [90] , CVDN contains dialogues between humans to guide the Agent to navigate in the environment, and the task design emphasizes cooperativeness, which is suitable for studying human-computer interaction and cooperative navigation, but the multiple rounds of dialogues and collaborative tasks increase the complexity of model training. the complexity of model training. The existing dataset has a large amount of data, but the diversity about objects is insufficient to meet human needs in specific situations [74].
In terms of modeling, the existing LLM navigation model utilizes the powerful natural language processing capabilities of LLM with extensive knowledge to guide the Agent in various environments. Since LLM is capable of understanding and generating natural language text, it can well simulate human language or accomplish tasks through conversations [55]. In addition, LLMs fine-tuned in various ways can handle multimodal data, such as combining multimodal inputs such as text, images, and speech, which can improve the accuracy of navigation and enhance the user interaction experience. However, LLM requires a large amount of high-quality data to train the model, and insufficient or inaccurate data will affect the navigation effect [62]. Although most of the existing LLM-based navigation models use pre-trained or directly invoked LLMs, they are not directly and task-specific trained, which can also adversely affect the final navigation results. Compared to non-LLM approaches, which do not need to rely on large amounts of natural language data, have relatively low data requirements, rely on image-text matching, and do not have the same scope of knowledge and comprehension as LLMs. LLMs have limitations in their ability to deal with fine-grained navigational instructions and spatial direction inference, and perform especially poorly in tasks that require long-term planning and spatial reasoning. In addition, LLMs are not sufficiently adapted to interactive navigation and show limitations in navigation tasks in environments containing traversable obstacles  [97]. Taken together LLM-based navigation methods are suitable for complex dialog and multimodal tasks, whereas non-LLM navigation methods  [57]perform well in image-text matching and visual tasks, with the exact choice depending on task requirements and resource conditions.
In models like NaVid [98], continuous RGB video frames serve as observational inputs, while instructions function as language inputs. Visual information from each frame is extracted using a frozen observation encoder, while instructions are transformed into language tokens using a text encoder. The information is then segmented into past frame data, current frame data, and navigation information, and Vicuna-7B is used to obtain the next action command. This multimodal fusion approach [99], along with the use of pre-trained models [100], is widely adopted in current LLM-based VLN models. The integration of multimodal information, such as in NaVid, including temporal, visual, and instruction information, can effectively enhance the model’s efficacy, while pre-training methods can alleviate training costs. This is also the primary direction for the future development of LLM-based VLN models.
When exploring the development and integration of LLMs in physically intelligent systems, it is necessary to consider the powerful environment-aware understanding provided by the development of LLMs, and the navigation performance of robots can be significantly improved by integrating LLMs with other modules, such as multi-sensor information fusion techniques. In addition, the predictive capabilities of LLMs can be further enhanced by introducing methods to predict future behaviors, enabling them to better understand and predict future events. For navigation tasks, studies have shown that the navigation ability of robots in complex environments can be effectively enhanced by integrating multimodal inputs such as vision, hearing, and haptics, and by utilizing complex decision-making frameworks [101, 102]. Future research should focus on developing generalized models that can handle multimodal inputs and perform complex tasks, while ensuring that these models remain efficient and flexible across different application scenarios. This includes, but is not limited to, the use of LLMs for multi-task learning and the development of new algorithms to optimize the robot’s behavioral strategies to adapt to changing environments [69]. According to the study, path planning can be performed more efficiently by using A* search algorithms and incorporating human feedback, while targeting in navigation specific behavioral control can be done by using LLMs to generate low-level control commands to directly control the robot’s dynamic motion without task-specific fine-tuning. This can be achieved by collecting cues from the physical environment, thus enabling the robot to perform basic actions such as walking [103]. At the same time, it is important for embodied subjects to form “memories” and to be able to avoid “forgetting” errors. LLMs may miss critical information when executing robot programs. Based on this, cue engineering strategies should be designed to minimize execution errors. By enhancing the multimodal processing capability of LLMs, using multimodal instruction tuning and multimodal learning strategies, the cross-modal dialog capability of LLMs can be greatly improved, and then they can be more capable of navigating tasks. In practice, multimodal command tuning technology can realize more accurate user intent understanding by combining voice commands and image input. By jointly learning data from different modalities such as speech, image and text, a richer and more comprehensive navigation model can be built [104].
In the interaction between embodied intelligence and the actual environment, the issues of data security and user privacy protection are particularly prominent. This is not only a technical issue, but also involves social ethics, legal norms and technological innovation. Meanwhile, the decision-making process of LLMs is often black-box, and it is difficult for developers to fully understand its internal mechanisms. In navigation applications, if users are unable to understand how LLMs make specific navigation recommendations, this may affect their trust and acceptance of the system. In terms of ensuring data security and protecting user privacy, a series of stringent measures must be taken to ensure that user data are not misused, leaked or tampered with. Some methods are used in protecting user privacy, such as training with synthetic data [105] or monitoring models’ outputs [106]. Therefore, privacy protection needs to be taken into account at the design stage, which means that it should be regarded as a fundamental principle rather than an afterthought remedy during the construction of embodied intelligence systems. The security of data collection, transmission, storage and processing should be considered at an early stage of system design, and appropriate encryption, access control and authentication mechanisms should be adopted to ensure the confidentiality and integrity of user data. In addition, a strict data access authority system should be established to limit the scope and purpose of data use to avoid data abuse or illegal use [107].

### v5

@@ III-C3 Comparison
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{L_{i}}{\max(P_{i},L_{i})}$
We have compared the works mentioned above, evaluating aspects such as the large models employed, application domains, and additional features. These comparisons are summarized in Table II.At the same time, we have divided the aforementioned works into two categories: Planning, as illustrated in fig.4, and Semantic Understanding, as shown in fig.5. Corresponding examples are provided for each category. Furthermore, we have analyzed the performance of these works across various datasets, the details of which are provided in Tables IV and V. It is important to clarify that comprehensive evaluation was unfeasible for some works due to the unavailability of public code. Consequently, the results are based solely on performance metrics reported in the corresponding publications. For an in-depth understanding of each work, readers are directed to the cited literature.
TABLE IV: Performance Comparison (Part I)
Benchmark Habitat RoboTHOR Gibson
SPL SR SPL SR SPL SR
CoW 6.3% - 10% 16.3% - -
ZSON - - - - 28% 36.9%
L3MVN - - - - 37.7% 76.1%
VLMAP 6% 2.5% - - - -
ESC - - 22.2% 38.1% - -
TABLE V: Performance Comparison (Part II)
Benchmark HM3D MP3D REVERIE R2R
SPL SR SPL SR SPL SR SPL SR
ZSON 12.6% 25.5% 4.8% 15.3% - - - -
CLIP-Nav - - - - 4.56% 15.2% - -
L3MVN - - 23.1% 50.4% - - - -
OVRL 29% 64% - - - - - -
ESC 22% 38.5% 13.7% 28.6% - - - -
NavGPT - - - - - - 29% 34%
BEVBert - - - - - - 64% 75%
MIC - - - - 41.97% 55.74% - -
SayNav 34% 60.32% - - - - - -
III-D Datasets
@@
a) MatterPort3D (MP3D) [18] is a comprehensive dataset containing 10,800 panoramic views from 194,400 RGB-D images across 90 architectural settings, focusing on 3D indoor environments. It includes RGB images, depth images, and semantic annotations for surfaces, camera poses, and segmentation, making it a key resource for self-supervised tasks and embodied AI. MP3D’s detailed scenes and global perspectives make it ideal for object detection, scene segmentation, and 3D reconstruction, although its focus on indoor environments limits its broader applicability.
b) TOUCHDOWN [22] is designed for navigation and spatial reasoning using Google Street View, containing 39,641 panoramic images from New York City. It features two tasks: Navigation and Spatial Description Resolution (SDR), emphasizing spatial reasoning and linguistic expertise. TOUCHDOWN is suitable for outdoor navigation tasks, unlike MP3D, which focuses on indoor settings, but the complexity of urban environments and inconsistent language annotations pose challenges.
c) Room-to-Room (R2R) [7] is a dataset for Vision-and-Language Navigation (VLN), extending MP3D with natural language instructions. R2R focuses on navigation in indoor spaces, requiring models to understand natural language and correlate it with visual information. This dataset presents a challenging research platform for complex decision-making and long-term memory in indoor navigation.
d) CVDN [120] is a dataset of over 2,000 ”human-to-human” dialogues in a photorealistic simulated domestic environment. It aims to train agents to navigate residential and commercial spaces using linguistic directives. Unlike R2R, CVDN features longer dialogues and emphasizes cooperative navigation, making it ideal for studying human-computer interaction and dynamic inquiry in navigation.
e) REVERIE [95] includes 10,567 panoramic images and 21,702 crowd-sourced instructions for target-oriented navigation tasks. It differs from traditional VLN datasets by providing semantically rich tasks, such as ”fold the towel in the bathroom with the fishing theme,” which requires long-term memory and semantic understanding. REVERIE is suitable for remote entity referencing and multi-task learning.
f) RxR [58] extends R2R with multilingual natural language instructions in English, German, and Spanish. It introduces spatiotemporal grounding, requiring agents to understand multilingual instructions, spatial reasoning, and long-term planning. RxR is a challenging dataset for advancing multilingual VLN.
g) ProcTHOR [34] is a dataset of 10,000 algorithmically generated 3D residential scenes designed for training embodied AI agents. It supports tasks like navigation, object manipulation, and scene understanding, and has shown robust performance across benchmarks such as Habitat 2022 and RoboTHOR challenges.
@@
TABLE VI: Comparison of Datasets. This table is a comparison among nine datasets, including seven evaluation methods.
@@
h) X-Embodiment[126] is an embodied robot manipulation dataset with over one million robot trajectories from 22 instances, covering 527 skills and 160,266 tasks. It merges data from 60 pre-existing robot datasets and standardizes them into the RLDS format[102], facilitating easy integration with various deep learning frameworks. This large-scale, diversified dataset is crucial for developing generalist robotic policies that can adapt to novel robots, tasks, and settings. By aggregating data across multiple platforms, X-Embodiment enables the creation of more versatile and universally applicable robotic policies, advancing the field of robot learning.
IV Challenges and Future Directions
LLMs face several challenges and limitations in navigation applications. Different datasets have unique strengths for various navigation and language guidance tasks. For instance, Matterport3D, R2R, RXR, and REVERIE [18][7][58][95] are suitable for indoor navigation tasks, while TOUCHDOWN [22] is tailored for urban environments, and ProcTHOR [34] caters to controlled environments. R3ED [175] is designed for long-term tasks with a memory module. MP3D provides high-quality RGB-D images and 3D reconstruction data but requires high hardware configurations [18]. TOUCHDOWN offers urban street scene data but may have inconsistent language annotations [100]. CVDN [120] emphasizes cooperative navigation, with longer dialogues and path descriptions, while ProcTHOR allows for customizable virtual environments, ideal for embodied AI tasks.
LLM-based navigation models leverage powerful NLP capabilities to guide agents across environments, capable of understanding and generating natural language [96][107]. However, these models typically rely on pre-trained LLMs without task-specific fine-tuning, affecting performance. Non-LLM methods, by contrast, require less data and avoid the need for extensive natural language understanding, excelling in tasks like image-text matching and visual tasks. LLMs, however, struggle with fine-grained navigation instructions and spatial reasoning, performing poorly in long-term planning and environments with traversable obstacles[154][14].
Models like NaVid [165] integrate multimodal data (RGB video, text) to enhance navigation, with pre-trained models alleviating training costs [163]. Future LLM-based models may benefit from multi-sensor fusion and prediction methods to improve navigation tasks, especially in complex environments [160][127]. The integration of multimodal instruction tuning can improve task performance, and privacy protection must be prioritized in the design of embodied AI systems [54] [164].